# Chemist-X: Large Language Model-empowered Agent for Reaction Condition Recommendation in Chemical Synthesis

### v5

@@ 2.1.3 Unit Tests for Phase Three
Phase Three develops a novel RCR algorithm with CL-SCL fingerprint as a proof of concept to show Chemist-X’s ability to utilize updated CAD tools. Unit tests in this subsection focus on the efficacy of the proposed CAD scheme, while a systematic validation of the agent is presented in the next subsection via web-lab experiments. To examine the generalization ability and the universality of the fingerprint, experiments are conducted as follows: we use a Buchwald-Hartwig reaction dataset [20] for model training and test the model with a Suzuki-Miyaura reaction dataset [28]. Further, in the evaluation, we follow the traditions established in prior RCR works [21] and use the maximum observed yield as the figure of merit. We denote the maximum observed yield of final recommendations by $\mu_{N}$, where $N$ is the batch size (BS) of the final recommendation.
@@
Benchmarking Experiments: We first compare our CL-SCL fingerprint against two reaction fingerprints from prior studies: DRFP [29] and Mordred [30]. We validate these fingerprints with three ML algorithms, i.e., random forest (RF) [31], tabular transformer (TT) [9], and Xgboost (XGB) [32], under different BS configurations. Table 1a presents the resulting $\mu_{N}$ across different experimental setups, which reveal that our CL-SCL fingerprint consistently delivers superior reaction condition recommendations than other alternatives regardless of BS setting and the ML model harnessed.
@@
(a) Experimental Results of Benchmarking Experiments ($\mu_{N}$ is recorded in the form of percentage)
@@
Table 1: A detailed result record for experiments conducted in Phase Three.
Ablation Study: There are two critical operations in generating the fingerprint: 1) the CIMG molecule encoder and 2) the CL-SCL network. A question one may ask is the effectiveness and necessity of each operation. To address this query, we conducted an ablation study that tests the indispensability of each element. The ablation study sets the BS as 5. In Experiment (i), we remove the CL-SCL network, allowing the chemical reactions encoded by CIMG to directly interface with subsequent ML models. In Experiment (ii), we substitute the CIMG encoder with a trivial encoder that simply concatenates molecule encodings. Beyond these alterations, the rest experimental setting mirrors the benchmarking experiment.
Table 1b presents the resulting $\mu_{5}$ of each experiment. As the table suggests, our original method has a significantly higher $\mu_{5}$ than case (i) and case (ii), regardless of the ML model applied. This finding indicates that both the CIMG encoder and the CL-SCL network are important components in Phase Three.
2.2 System Validation via Wet-lab Experiments
@@ 2.2.1 Experimental Setup
To obtain 6-(1-methyl-1H-indazol-4-yl), the target product, we consider the following chemical subspace encompassing four key dimensions: reactant, ligand, base, and solvent. Further, considering the typical reaction conditions of Suzuki reactions and the availability of chemicals, the initial chemical subspace, as shown in Figure 4a and 4b, has 10000 possible reactions within it. [^1] The chemical subspace consists of nine reactants (1a-1c react with 2c-2d and 1d-1e react with 2a-2b, leading to 10 combinations), ten ligands, ten bases, and ten solvents, which leads to in 10000 possible reactions.
Figure 4: Overview and procedure of the studied Suzuki-Miyaura reactions. a) The chemical space of the studied Suzuki-Miyaura reaction. b) Structure of phosphorus ligands shown in a). c) All reactions were conducted in Schlenk tube under Ar atmosphere. d) Identification of the product and yield analysis with LC-MS and HPLC. e) The Calculation of the Relative Response Factor of internal standard method; f) Yield analysis of each reaction with HPLC.
After operation and analysis conducted in Phase One and Phase Two, our agent identifies from existing literature that K2CO3 could be the optimal base, and dioxane/H2O could be the most suitable solvent. This insight effectively reduces the unknown variables within our chemical subspace to two dimensions: reactants and ligands. Consequently, the reaction subspace is simplified to 100 viable reaction configurations. Then, Phase Three, using the TT network as the example ML model embedded after the CL-SCL network, makes final recommendations among the narrowed reaction subspace. In a practical application, Phase Three is used only once with fixed CL-SCL and TT networks. However, to comprehensively test our scheme’s robustness in the face of randomness during the training process, we re-examine Phase Three three times, with each time using a different random seed in the training process.
@@ 2.2.2 Experimental Details and Discussions
We now give a detailed description of our operation in the experiments. Unless otherwise specified, all reagents and solvents were commercially available and used without further purification. All reactions were performed under a positive pressure of argon atmosphere and preliminary monitored by thin-layer chromatography (TLC) while visualized with ultraviolet (UV) light at 254 nm. The product was purified by column chromatography on silica gel 200–300 mesh and identified by high-performance liquid chromatography/mass spectrometry (HPLC-MS), just as depicted in Figure 4d. Fourth, the reaction yield used 4-chloro-6,7-dimethoxyquinoline as an internal standard and was analyzed by HPLC, which was determined by Thermo Scientific Vanquish Core HPLC with a binary pump, a UV detector, an Acclaim 120-C18 (3*150 mm, 3 $\mu$m), using mobile phase A: H2O /B: CH3CN, the flow rate is 0.5 mL/min.
A typical procedure for the synthesis process and yield analysis is as follows: A 20 mL sealed Schlenk tube under argon atmosphere was charged with 6-bromoquinoline (1b, 129 mg, 0.5 mmol), 1-methyl-4-(4,4,5,5-tetramethyl-1,3,2-dioxaborolan-2-yl)-1H-indazole (2d, 103 mg, 0.5 mmol), Pd(OAc)2 (5.66 mg, 0.025 mmol), potassium carbonate (B1, 173 mg, 1.25 mmol) and Amphos (L10, 13.3 mg, 0.05 mmol) in dioxane/H2O (9:1, 5 mL), as shown in Figure 4c. The reaction mixture was heated at 85oC for 16 hours, cooled down to room temperature, and monitored by TLC. Then reaction solution was diluted to 10 mL by methanol, measuring off 0.5 mL diluted solution by pipette, added in 5 mg internal standard, and re-diluted to 10 mL, which was used for HPLC analysis. As the Relative Response Factor (RRF=0.5343) has been determined in Figure 4e and the peak area ratio of product to internal standard is determined in Figure 4f, it is sufficient to calculate the yield is 91.6%. Other reactions were conducted and analyzed following procedures similar to the above steps.
@@ 3 Method
3.1 Phase One: information retrieval using molecule database memory
Databases like PubChem typically offer API interfaces for user queries, along with comprehensive tutorials that demonstrate the functions of these APIs through examples. In Phase One, we harness the API and its documentation with our innovative TMS-ICL strategy.
@@ 3.1 Phase One: information retrieval using molecule database memory
3.2 Phase Two: information analysis using online literature memory
Phase Two encompasses two subtasks. The first involves creating web crawlers that extract HTML content from scientific literature websites. We employ web crawlers for data retrieval in this phase, as literature platforms like ChemSpider offer their resources through web interfaces, unlike the APIs utilized in Phase One. The second subtask is to analyze the obtained HTML data and generate a list of potential reaction conditions identified according to the studied literature, which serves as Phase Two’s deliverable.
The development and operation of the web crawlers are guided by ICL prompts. To ensure precise and efficient HTML data extraction, we have devised a specialized prompting scheme for our crawler-based information retrieval. The protocol unfolds as follows. The initial step instructs the agent on setting up the crawling environment, including the installation of necessary Python packages, opening the browser, and logging in to the literature platform with the given username and password. The second step outlines the structure of the target webpage, prompting the agent to emulate human interaction (i.e., clicking and typing) using the crawler. Once the interactions are complete, the third step guides the agent to download the HTML content featuring the search results for the molecule of interest. The final prompt asks the crawler to iterate over the third and fourth steps until all similar molecules identified in Phase One have been searched within literature platforms, with all relevant HTML sources downloaded. Fig. 2 of the extended data shows an example Python realization for an interested molecule.
After downloading these HTML sources, our agent is tasked with composing an HTML analysis function in Python. The example we design for the ICL process should consider complex HTML structures (e.g., multiple reactions are found for the target molecule, and each reaction may consist of multiple possible conditions reported in previous works) to ensure the generated code’s effectiveness under intricate circumstances. This approach contrasts with general HTML analysis methods [17, 18], in which the raw HTML source is directly given to the LLM. Our method is explicitly tailored for the webpages of literature platforms and allows for code reuse in subsequent HTML file analysis. Experimental results in Section 2.1.2 have demonstrated that our approach not only improves the performance of HTML analysis but also minimizes computational resources consumed. We refer readers to Fig. 3 of the extended data for an example code realization completed by GPT-4.
3.3 Phase Three: final recommendation with pre-packaged fingerprint tool
It is important to note that a reaction condition emerging from Phase Two may be incomplete. Although a literature platform could offer comprehensive condition information for numerous chemical reactions, it might inadvertently omit critical details in a reaction, such as reagents, catalysts, solvents, and yield. In other words, detailed reaction data from the original paper might be uncaptured in the literature platform. Chemist-X aims to accelerate chemical reaction designs and facilitate future robot-operated chemical reaction platforms that free human labor by automatically providing complete recommendations with no information loss. In light of this, at the beginning of Phase Three, we fill in the overlooked information with a comprehensive set of potential solutions in our chemical subspace.
After the information-filling process, the agent needs to identify the best possible reaction conditions with the highest expected yield. Previous works have proved that advanced ML algorithms are competent for the RCR task [20, 19, 21, 34]. However, one major problem we noticed in these works is that their reaction encodings, also known as reaction fingerprints, are just straightforward concatenations of encoded molecules. Although these reaction encoding schemes are easy to implement, we want to emphasize the importance of having a reaction fingerprint with additional yield information integrated into it when training an ML model for yield-predicting tasks or RCR tasks targeting optimal yields.
@@ 3.3.1 Reaction data processing
In this paper, we tested three different ML models: random forest [31], tabular transformer [9], and Xgboost [32]. Implementations of the three ML models are packaged within Phase Three’s API function. When calling the API, Chemist-X can select one of the three models for the execution of Phase Three via a user-friendly API interface.
@@ 4 Conclusion
This paper introduces Chemist-X, an innovative RAG-AI agent that harnesses the latest research breakthroughs in RAG technology and LLMs to transform the landscape of reaction condition recommendations in chemical synthesis. In general, this paper makes three significant contributions. First, with the “search-analyze-recommend” pattern of expert chemists and the novel reaction fingerprint that integrates yield data, Chemist-X demonstrates remarkable proficiency in automating the RCR task. Second, the use of updated online databases as knowledge sources allows our agent to transcend the limitations of traditional AI systems, which are confined to the knowledge within their training data. Furthermore, the agent’s ability to leverage cutting-edge CAD tools also enables the continuous evolution of the system. Third, the successful application of our RAG-AI agent to the Suzuki reaction example, validated through wet-lab experiments, underscores the practical impact of this research. With the above technical and experimental contributions, this research brings the vision of fully automated chemical discovery one step closer, marking a significant leap in the field of computer-aided chemistry.

### v6

@@ 2.2 Phase Two Unit Tests
Phase Two develops a novel RCO algorithm with CL-SCL fingerprint, as a proof of concept, to show Chemist-X’s ability to utilize updated CAD tools. Unit tests in this subsection focus on the efficacy of the proposed CAD scheme, while a systematic validation of the agent is presented in the next subsection via wet-lab experiments. To examine the generalization ability and the universality of the fingerprint, experiments are conducted as follows: we use a Buchwald-Hartwig reaction dataset [15] for model training and test the model with a Suzuki-Miyaura reaction dataset [26]. Further, in the evaluation, we follow the traditions established in prior RCO works [16] and use the maximum observed yield as the figure of merit. We denote the maximum observed yield of final recommendations by $\mu_{N}$, where $N$ is the batch size BS of the final recommendation.
@@
Benchmarking Experiments: We first compare our CL-SCL fingerprint against two reaction fingerprints from prior studies: DRFP [27] and Mordred [28]. We validate these fingerprints with three ML algorithms, i.e., random forest (RF) [29], tabular transformer (TT) [30], and Xgboost (XGB) [31], under different batch size configurations. Table 1a presents the resulting $\mu_{N}$ across different experimental setups, which reveals that our CL-SCL fingerprint consistently delivers superior reaction condition recommendations than other alternatives regardless of batch size setting and the ML model harnessed.
@@
(a) Benchmarking Results ($\mu_{N}$ is recorded in the form of percentage, and BS refers to batch size)
@@
Table 1: A detailed result record for experiments conducted in Phase Two.
Ablation Study: As presented in the methodology section in detail, there are two critical operations in generating the fingerprint: 1) the CIMG molecule encoder and 2) the CL-SCL network. A question we are interested in is the effectiveness and necessity of each operation. To address this query, we conducted an ablation study that tests the indispensability of each element. The ablation study sets the batch size as 5. In Experiment (i), we remove the CL-SCL network, allowing the chemical reactions encoded by CIMG to directly interface with subsequent ML models. In Experiment (ii), we substitute the CIMG encoder with a trivial encoder that simply concatenates molecule encodings. Beyond these alterations, the rest of the experimental setting mirrors the benchmarking experiment.
Table 1b presents the resulting $\mu_{5}$ of each experiment. As the table suggests, our original method has a significantly higher $\mu_{5}$ than case (i) and case (ii), regardless of the ML model applied. This finding indicates that both the CIMG encoder and the CL-SCL network are important components in Phase Two.
2.3 Phase Three Wet-Lab Validations
This subsection presents experimental details about the agent-controlled web-lab experiments. We will start with our experimental setups on the IC platform, and then we will present experimental results and discuss them accordingly. For technical details on how the control is realized, we refer readers to Section 3, the methodology section, for details.
@@ 2.3.1 Experimental Setup
To obtain 6-(1-methyl-1H-indazol-4-yl), the target product, we consider the following chemical subspace encompassing four key dimensions: reactant, ligand, base, and solvent. Further, considering the typical reaction conditions of Suzuki reactions and the availability of chemicals, the initial chemical subspace, as shown in Figure 3a and 3b, has 10,000 possible reactions within it. [^1] The chemical subspace consists of nine reactants (1a-1c react with 2c-2d and 1d-1e react with 2a-2b, leading to 10 combinations), ten ligands, ten bases, and ten solvents, which leads to in 10000 possible reactions.
Figure 3: Overview and procedure of the studied Suzuki-Miyaura reactions. a) The chemical space of the studied Suzuki-Miyaura reaction. b) Structure of phosphorus ligands shown in a). c) Unchained Lab in the IC platform, equipped with modules for storage of solid, liquid reagents and solvents, pipettes for viscous or trace liquid reagents, vials for different experiments, an analytical balance and reactors (OSR). d) Reactors (OSR) of the Unchained Lab with inert gas protection. e) The orbital robot and on-island robotic arm of the IC platform for achieving automated transfer operations. f) Identification of the product and yield analysis with the HPLC-MS and HPLC systems. g) The calculation of the Relative Response Factor with the internal standard method. h) Yield analysis of each reaction with HPLC.
After operation and analysis conducted in Phase One, our agent identifies from existing literature that K2CO3 could be the optimal base, and dioxane/H2O could be the most suitable solvent. This insight effectively reduces the unknown variables within our chemical subspace to two dimensions: reactants and ligands. Consequently, the reaction subspace is simplified to 100 viable reaction configurations. Then, Phase Two, using the TT network as the example ML model embedded after the CL-SCL network, makes final recommendations among the narrowed reaction subspace. In a practical application, Phase Two is used only once with fixed CL-SCL and TT networks. However, to comprehensively test our scheme’s robustness in the face of randomness during the training process, we re-examine Phase Two three times, with each time using a different random seed in the training process.
@@ 2.3.2 Experimental Details and Discussions
We now give a detailed description of our operation in the experiments. Unless otherwise specified, all reagents and solvents were commercially available and used without further purification. All reactions were performed in the optimization sampling reactor (OSR) of the Unchained Lab, which enables synthesis reactions requiring conditions of heating, pressurization, continuous stirring and inert gas protection (3c). The product was purified by column chromatography on silica gel 200–300 mesh and identified by high-performance liquid chromatography/mass spectrometry (HPLC-MS). Fourth, the reaction yield used 4-chloro-6,7-dimethoxyquinoline as an internal standard and was analyzed by HPLC, which was determined by Thermo Scientific Vanquish Core HPLC with a binary pump, a UV detector, an Acclaim 120-C18 (3*150 mm, 3 $\mu$m), using mobile phase A: H2O /B: CH3CN, the flow rate is 0.5 mL/min.
A typical procedure for the synthesis process and yield analysis is as follows: Before synthesis, the solid, liquid reagents and solvents were pre-prepared in storage vials, and the air in OSR was replaced with inert gas in advance. The experiment operations of the robotic system for Suzuki-Miyaura coupling, including pre-treatment, synthesis, post-treatment and characterization of the reaction mixture, were automatically performed using the Unchained Lab and robotic arms of the IC platform without human intervention (Supplementary movie 1). First, the reaction mixture was prepared by automatically adding substrates, catalyst, ligand, base and solvent into a 20 mL vial for pre-treatment. For example, for performing the first condition experiment of the first batch in Table 2, 6-bromoquinoline (1b, 312 mg, 1.5 mmol), 1-methyl-4-(4,4,5,5-tetramethyl-1,3,2-dioxaborolan-2-yl)-1H-indazole (2d, 387 mg, 1.5 mmol), Pd(OAc)2 (16.9 mg, 0.075 mmol), potassium carbonate (B1, 519 mg, 3.75 mmol) and Amphos (L10, 39.8 mg, 0.15 mmol) in dioxane/H2O (9: 1, 15 mL) were added into a 20 mL vial. Five reactions with different conditions were performed in parallel for each batch of experiments (Table 2). Then, the reaction mixture solutions were transferred to the OSR (Figure 3d), heated at 85oC with continuous stirring for 16 h under nitrogen atmosphere, then cooled down to room temperature. For post-treatment, 5 mg of the internal standard (IS) was dissolved in 10 mL of methanol (MeOH), and then 100 $\mu$L of the reacted mixture was aspirated and added to the IS stock solution, shaken until thoroughly mixed and transferred to 2 mL vials to prepare the mixture solution for characterization. Then, through the cooperation of an orbital robot and an on-island robotic arm (Figure 3e), the vial plate, containing a batch of 5 mixture solutions obtained under different conditions, was transferred to the HPLC island for characterization (Figure 3f).
With the IS and the reference standard (i.e., product purified by column chromatography), the relative response factor (RRF) was determined with the following equation:
$\text{RRF}=\frac{{{A}_{\text{IS}}}\times{{C}_{\text{reference}}}}{{{A}_{\text{reference}}}\times{{C}_{\text{IS}}}}$ (1)
where $A_{\text{IS}}$ and $A_{\text{reference}}$ are the peak areas of the IS and the reference standard obtained by the HPLC system (Figure 3g), $C_{\text{reference}}$ and $C_{\text{IS}}$ are the known concentrations of the reference standard and the IS.
For the quantitative characterization of the target product in the reaction mixture solution, $C_{\text{product}}$ was calculated with the equation:
${{C}_{\text{product}}}=\frac{{{A}_{\text{product}}}\times{{C}_{\text{IS}}}\times\text{RRF}}{{{A}_{\text{IS}}}}$ (2)
where $C_{\text{product}}$ is the concentration of the product in the reacted mixture solution to be determined, $A_{\text{product}}$ and $A_{\text{IS}}$ are the peak areas of the product and the IS obtained by the HPLC system (Figure 3h), $C_{\text{IS}}$ is the known concentration of the IS in the reacted mixture solution, RRF is calculated by equation 1.
Then the product yield of the reaction was quantified with the equation:
[…]
where $C_{\text{substrate}}$ is the known concentration of the substrate added in the unreacted mixture solution. The experiments at other reactions were conducted and their mixture solutions were analyzed following procedures similar to the above steps.
@@ 3 Method
3.1 Phase One: Information Retrieval Using Molecule and Literature Databases
Databases like PubChem and Amazon EC2 typically offer API interfaces for user queries, along with comprehensive tutorials that demonstrate the functions of these APIs through examples. In Phase One, we harness the API and its documentation with our innovative TMS-ICL strategy.
@@ 3.1 Phase One: Information Retrieval Using Molecule and Literature Databases
3.2 Phase Two: Reaction Condition Optimization with Pre-packaged Fingerprint Tool
It is important to note that a reaction condition emerging from Phase One may be incomplete. Although a literature platform could offer comprehensive condition information for numerous chemical reactions, it might inadvertently omit critical details in a reaction, such as reagents, catalysts, solvents, and yield. In other words, detailed reaction data from the original paper might be uncaptured in the literature platform. Chemist-X aims to accelerate chemical reaction designs and facilitate future robot-operated chemical reaction platforms that free human labor by automatically providing complete recommendations with no information loss. In light of this, at the beginning of Phase Two, we fill in the overlooked information with a comprehensive set of potential solutions in our chemical subspace.
After the information-filling process, the agent needs to identify the best possible reaction conditions with the highest expected yield. Previous works have proved that advanced ML algorithms are competent for the RCO task [15, 14, 16, 33]. However, one major problem we noticed in these works is that their reaction encodings, also known as reaction fingerprints, are just straightforward concatenations of encoded molecules. Although these reaction encoding schemes are easy to implement, we want to emphasize the importance of having a reaction fingerprint with additional yield information integrated into it when training an ML model for yield-predicting tasks or RCO tasks targeting optimal yields.
@@ 3.2.2 The CL-SCL Network
In this paper, we tested three different ML models: random forest [29], tabular transformer [30], and Xgboost [31]. Implementations of the three ML models are packaged within Phase Two’s API function. When calling the API, Chemist-X can select one of the three models for the execution of Phase Two via a user-friendly API interface.
3.3 Phase Three: Automated Laboratory Execution with LLM-Supervised Interface Control
Many laboratory automation platforms lack user-accessible APIs for machine control, often relying on graphical user interfaces (GUIs) that require manual interactions such as mouse clicks and keyboard inputs [37, 38, 39]. To address this challenge, Chemist-X employs a novel LLM-supervised interface control framework that enables the automated execution of chemical experiments by replicating human interactions with laboratory software GUIs. [^2] For a detailed Python implementation of our equipment control framework, we refer readers to our GitHub repository [40].
To facilitate automated execution, Chemist-X is provided with a demonstration video that records a human operator interacting with the experimental software. This video serves as a reference for the LLM-based agent to understand and replicate the procedural workflow. The recorded demonstration typically consists of three key stages, including 1) Initialization: Predefined actions such as launching the experiment control software and ensuring all necessary modules are loaded; 2) Reaction Condition Configuration: Selection of appropriate reactants, solvents, catalysts, and substrates, along with specification of experimental parameters (e.g., temperature, reaction duration); and 3) Execution: Verification of the experiment setup and initiation of the reaction.
Among these stages, the initialization stage and the execution stage follow a fixed set of operations and can be automated by tracking and replicating mouse movements and click behaviors. In contrast, reaction condition configuration presents a greater challenge due to its dependency on specific reaction requirements. Figure 4a provides a detailed illustration of how our novel multi-step prompt framework enables the LLM to understand, analyze, and replicate human interactions.
To begin with, we provide the LLM with the demonstration video alongside a general task description to ensure the model’s comprehensive understanding of the workflow. A key frame detection algorithm is then applied in Step Two to filter out irrelevant frames (e.g., idle periods of the recorded video, during which time no human operation is conducted). For the detected key frames, we then apply a clustering algorithm to identify the mouse pointer locations to capture the user’s clicking operations. The extracted information about the mouse’s location is then passed to the LLM, which generates simulated click sequences (based on the pyautogui Python package) for the fixed operation in Initialization and Execution stages. Note that engineering details of the key frame detecting algorithm and the mouse identification algorithm are not presented here for better scientific focus. Interested readers can refer to our open-source implementation in the GitHub repository [40].
Figure 4: Technical details about the LLM-supervised control script generation for wet-lab equipments. Figure a) presents our prompting framework. Figure b) gives the ablation study result for the mouse clicking algorithm. Figure c) shows the parameter sensitivity of the LLM-supervised system.
Step Three handles the more complex Reaction Condition Configuration stage. Unlike the Initialization and Execution stages, which involve fixed operations, this stage requires adaptive decision-making based on the specific experiment being executed. The demonstration video only provides an example of human interaction, but the exact chemical parameters and selection choices may vary across different experiments. Therefore, directly mimicking recorded mouse clicks is insufficient. To address this, Chemist-X uses an optical character recognition (OCR) tool to extract textual information (e.g., reagent names, concentrations, dropdown menu items) from the software interface and then interprets the extracted text and matches this textual information with the desired reaction conditions (see the third prompt in Fig.4a for details). Following the matching results, the agent generates the associated control script (also based on the pyautogui Python package) that can execute mouse click sequences to configure the reaction setup as required.
Finally, Chemist-X combines the code generated in Step Two (i.e., for the Initialization and Execution stages) and Step Three (for the Reaction Condition Configuration stage) to obtain the desired control script and double-checks the code to fix potential errors.
@@ 4 Conclusion
This paper introduces Chemist-X, an innovative RAG-AI agent that harnesses the latest research breakthroughs in RAG technology and LLMs to transform the landscape of reaction condition recommendations in chemical synthesis. In general, this paper makes three significant contributions. First, with the “search-analyze-recommend” pattern of expert chemists and the novel reaction fingerprint that integrates yield data, Chemist-X demonstrates remarkable proficiency in automating the RCO task. Second, the use of updated online databases as knowledge sources allows our agent to transcend the limitations of traditional AI systems, which are confined to the knowledge within their training data. Furthermore, the agent’s ability to leverage cutting-edge CAD tools also enables the continuous evolution of the system. Third, the successful application of our RAG-AI agent to the Suzuki reaction example, validated through wet-lab experiments, underscores the practical impact of this research. With the above technical and experimental contributions, this research brings the vision of fully automated chemical discovery one step closer, marking a significant leap in the field of computer-aided chemistry.